# A Single Motor Nano Aerial Vehicle with Novel Peer-to-Peer Communication and Sensing Mechanism

### v1

@@ Abstract
Communication and position sensing are among the most important capabilities for swarm robots to interact with their peers and perform tasks collaboratively. However, the hardware required to facilitate communication and position sensing is often too complicated, expensive, and bulky to be carried on swarm robots. Here we present Maneuverable Piccolissimo 3 (MP3), a minimalist, single motor drone capable of executing inter-robot communication via infrared light and triangulation-based sensing of relative bearing, distance, and elevation using message arrival time. Thanks to its novel design, MP3 can communicate with peers and localize itself using simple components, keeping its size and mass small and making it inherently safe for human interaction. Here we present the hardware and software design of MP3 and demonstrate its capability to localize itself, fly stably and maneuver in the environment using peer-to-peer communication and sensing.
@@ I Introduction
Additionally, a wide variety of tasks, including search and rescue, transportation or drone shows, can involve direct or indirect interaction with humans. But highly capable multi-motor drones tend to be heavy and unsafe for human interaction without safety equipment [21], a risk that is amplified when interacting with a large number of UAVs. While there are impressive works in active avoidance of obstacles which partially addressed this issue [22, 17], it would be preferred if the drone could be light enough to be inherently safe [23] to interact with.
@@
In this paper, we present a single motor nano [30, 14] aerial vehicle, MP3, with a novel, minimalist peer-to-peer communication and sensing system. MP3 is similar to its predecessors: its only motor is offset from its center of mass to provide maneuvering ability similar to the original Maneuverable Piccolissimo (MP) [27] and its controller is similar to MP2 [15]. However, MP3 is equipped with a novel communication and sensing system as well as a completely new chassis, electrical and software design. MP3 is capable of communicating with neighboring beacons or other MP3s using infrared light (IR), and sensing its relative bearing, distance and elevation to neighbors using a single-drone triangulation method. Using this information, MP3 can subsequently determine its global Cartesian position to millimeter accuracy. Using the global position information, MP3 can control its position in 3D space with a single actuator. The simplicity of MP3 allows it to be 20.0g in weight and 112mm in diameter. The small mass and the partially enclosed propeller make MP3 inherently safe for human interaction even with collisions [23]. We present the hardware and software design of MP3, with emphasis on the novel communication and sensing system, and validate its capability with experiments.
Figure 2: An exploded view of MP3. MP3 has 7 major parts: 3D printed main and top chassis, a main PCB, a top transmitter, a motor-propeller assembly, and two batteries. Bottom transmitter is placed at the center of the bottom side of the main PCB.
@@ II-A Overall Concept
MP3, as shown in Fig. 1 and 2, is a minimalist single-rotor drone with on-board communication and position sensing capabilities using a novel optical communication and sensing system. Unlike ordinary drones, MP3 has only one motor offset from its center of mass, two rigid bodies, (rotor fixed with a propeller and stator fixed to the main chassis), and no inertial measurement unit (IMU). As the motor spins, the body of MP3 spins like a frisbee in the opposite direction as shown in Fig. 1, and the gyroscopic effect of the spinning body keeps the MP3 stable. By pulsing the thrust at different times, we can tilt the MP3 and control its 3D position in flight [15].
@@
Figure 4: MP3’s communication and sensing system. (a) shows the receiver system on MP3, (b) shows the top and bottom transmitter system (TX) on MP3, and (c) shows the transmitter system on the Beacon. Note that the transmitter system on the MP3 and on the Beacon are the same.
@@ II-B The Communication and Sensing System
Based on this principle, we designed the communication and sensing system hardware as shown in Fig. 4. Omindirectional transmitters built on MP3s blink their infrared (IR) LED to broadcast messages in the form of encoded pulses, and MP3s can receive messages and localize themselves using the three photodiodes built on them.
@@ II-B1 Theory of Operation
To illustrate how MP3s achieve relative position sensing and absolute localization, we consider a simplified theoretical model as shown in Fig. 5, where two point-like transmitters are constantly emitting light omnidirectionally and broadcasting their global position $\vec{p}_{1},\vec{p}_{2}$ via IR communication. A simplified MP3 is trying to determine its relative bearing, distance and elevation to the transmitters as well as its global position. The MP3 is modeled as a solid body rotating at a constant angular velocity $\omega\hat{z}$, and is equipped with three light receivers with planer fields of view (FOV), i.e. a receiver will only pickup light signal when the transmitter is in its FOV. Two light receivers (right and left) are oriented orthogonal to the MP3 horizontal plane, and one light receiver (middle) is oriented at an angle ($\phi$) from orthogonal. As the MP3 rotates, its left, middle, and right receiver will pickup signals from transmitter 1 at time $t_{\text{1L}},t_{\text{1M}},t_{\text{1R}}$ respectively, so we can determine the distance and elevation angle between the MP3 and transmitter 1 as:
@@
We can obtain $r_{2},\alpha_{2}$ from transmitter 2 similarly. Furthermore, at times $t_{1}=(t_{\text{1L}}+t_{\text{1R}})/2$ and $t_{2}=(t_{2L}+t_{2R})/2$, the MP3 is facing transmitter 1 and 2 respectively, so we can express the angle between transmitter 1 and transmitter 2 as
@@
While $\sigma_{\alpha_{1}}$ and $\sigma_{\theta_{12}}$ are on the same order of magnitude as $\omega\sigma_{t}$ and are not very sensitive to relative position between robots, $\sigma_{r_{1}}$ is proportional to $r_{1}^{2}$, meaning that the localization is more accurate when drones are closer to each other, which is helpful for collision avoidance in dense swarms.
Unfortunately, non-ideal factors make the implementation of this theory more complicated. These include non-ideal sensor optics, limited time resolution constraints by the communications protocols, etc.
@@ II-B3 Electrical Design
MP3 can only receive a communicated position when its receivers are facing a transmitting MP3. MP3s typically spin at 25Hz and each of their receivers have a ±$\pm 10^{\circ}$ horizontal FOV, meaning that, in each revolution, a MP3 only has approximately 2-5ms to communicate with another MP3, and in the worst case scenario this gives a communication latency of 40ms. To facilitate communication with multiple MP3s using time division multiple access (TDMA), each MP3 could take no more than 500~1000us to transmit all necessary information to a receiver. If we want to communicate tens of bytes over this period, we need a system capable of communicating at around 1Mbps. This means that we could not use commercially available IR transceivers but had to implement our own.
@@
The receiver circuit is shown in Fig. 8. PCB layout and filter are designed to reduce electromagnetic interference of the motor and filter out almost all environmental light. The analog part of the circuit has a gain of 67mV/uA, an input referred noise as low as 36nA RMS, and a -3dB passband of [0.42MHz,10.5MHz]. The overall circuit can reliably pickup photocurrent as low as 140nA, enabling all manufactured MP3s to communicate across at least 50cm free space with more than 95% success using non-directional transmission and reception.
@@ II-B4 Communication Protocol Design
When designing the communication protocol, we want it to be general purpose, to be simple to implement, to have multiplexing ability, and to provide accurate enough timing information $t_{\text{1L}},t_{\text{1M}},t_{\text{1R}}$ to facilitate position sensing.
To make our protocol general purpose, we want it to handle messages of different length and content in the same way. This would allow a MP3 to receive communicated positions from other MP3s, as well as other task-dependent messages in future work. A natural approach would be to use packets: a complete message is transmitted in one or multiple 30-bit packets, each containing 2 bytes of data and additional metadata like the ID of the MP3 transmitting the message, the ID of packet, the cyclic redundancy check (CRC) value of the message, etc. A MP3 can obtain timing information by looking at when it starts to receive packets from another MP3.
@@ II-B5 Global Localization Algorithm
We can now bring everything together. In section II-B1, we have already presented the basic theory to use the time when we received messages from other transmitters to determine relative position. The optical and electrical designs (section II-B2 and section II-B3) make MP3s as close to the theoretical model as possible. Finally, we can use the proposed communication system (section II-B) to obtain the global position of neighboring MP3s. In this section, we present a simple way of using these relative distance and angle measurements and neighboring MP3s’ global position information to determine MP3’s global position. We execute localization in three steps to simplify the computation without sacrificing much accuracy:
@@
    Determine the time $t_{x}$ when MP3 is facing $\hat{x}^{+}$ direction and global horizontal position $s_{X},s_{Y}$ using angular velocity of the drone $\omega$, the global horizontal position of the transmitters $(\vec{p}_{i})_{XY}$, the time when MP3 is facing these transmitters $t_{i}$, and the measured relative horizontal distance $r_{i}$.
@@
In the following discussion, we make the following assumptions:
@@
    MP3’s geometric axis is always vertical and it spins along its geometric axis.
@@
The first assumption will introduce error into vertical position estimation as MP3s’ geometric axis and angular velocity axis will deviate slightly, but usually they are only off by a few degrees and we are willing to sacrifice the precision in elevation for faster computation. The second and third assumptions are quite accurate as in experiments the angular velocity changes by <0.5% and the position usually changes by <5mm in one revolution. While the measured variables are not strictly Gaussian distributed, the fourth assumption is added for ease of explanation without compromising the core idea.
@@
For the second step, we will determine the most likely $t_{x},s_{X},s_{Y}$ given $\vec{p}_{i},r_{i},t_{i}$ and their uncertainty estimates $\sigma_{r_{i}},\sigma_{t_{i}}$.
$\mathop{\mathrm{argmin}}\limits_{t_{x},s_{X},s_{Y}}\sum\limits_{i=1}^{N}\left(\frac{\Delta_{X}^{2}+\Delta_{Y}^{2}}{\sigma_{r_{i}}^{2}}\right.\left.+\frac{\mathop{\mathrm{atan}}(\Delta_{X},\Delta_{Y})-\omega(t_{i}-t_{x})}{\omega^{2}\sigma_{t_{i}}^{2}}\right)$ (8)
@@
which is the root mean square (RMS) of uncertainty in X and Y direction. For simplicity of computation, in the third step we will assume $s_{X},s_{Y}$ are independent and both have uncertainty of $\sigma_{XY}$.
@@
where $\delta_{i}$ is the uncertainty of $(\vec{p}_{i})_{X}-\sqrt{\Delta_{X}^{2}+\Delta_{Y}^{2}}\tan\alpha_{i}$
@@
In experiments, we additionally applied an exponential filter with time constant of 0.06s to the position obtained ${s_{X},s_{Y},s_{Z}}$ to reduce the noise of position estimates.
@@ II-C The Mechanical System
Several other changes include: a brushless motor instead of a brushed motor to improve longevity and battery life, carbon fiber-infused nylon for chassis to reduce weight and increase strength, and wings added to the side of MP3 to improve passive stability and slow rotation rates.
@@ III Drone Flight Experiments
For each experiment, we collected the reference trajectory of the MP3s using two cameras, one over the apparatus and one on the side, and we also recorded the estimated trajectory of the MP3 which is computed on-board using only relative measurements and data communicated though our IR communication system. Data are recorded after MP3 is flying steadily, about 10~20s after launch.
@@ III-A Position Holding Experiment
In the position holding experiments, MP3 localizes itself by sensing three neighboring beacons. MP3 then tries to remain at a predefined set point at the center of the three beacons. The setup is shown in Fig. 10. We plot the reference trajectories of all flights in Fig. 11. The trajectories indicate successful and stable position holding, since MP3 is staying within 30mm around the origin most of the time (in contrast, the radius of MP3 is 56mm). The root mean square (RMS) of the drone’s reference position is {17.6, 22.5, 12.7}mm.
@@
Figure 12: Distribution and spectrum of localization error of MP3 in position holding experiments. (a) shows the localization error’s distribution and (b) shows the localization error’s spectrum.
A way to evaluate the precision of localization is to calculate the error between the robot-sensed trajectory and reference trajectory. Fig. 12(a) shows the distribution of the localization error, which resembles a normal distribution with $\sigma=\{2.4,2.0,3.6\}$mm respectively. This suggests that our sensing system and localization algorithm can achieve millimeter level accuracy. The distribution’s offset in X and Y direction may be due to inaccurate calibration of the camera over the entire flight arena or slight displacement of the beacons, and the offset in Z direction may also be caused by the MP3’s tilt.
Besides the localization error’s distribution, we are also interested in the spectrum of error, as the higher frequency components in error will be amplified more when we take derivatives to compute the velocity and acceleration. We split the obtained trajectory data into chunks of 5 seconds, and resample them at 100Hz. We then compute the discrete Fourier transform (DFT) of the error samples and plot the RMS density of error signal at different frequency in Fig. 12(b). The lower frequency (<2Hz) components are dominated by systematic error like the inaccurate placement of beacons or inaccuracy in the functions that converts time $t_{1L},t_{1M},t_{1R}$ to distance and elevation angle. These sources are dependent on the position of the drone which is slowly changing. The higher frequency (>2Hz) components, however, mainly come from the discrete and random nature of timing measurements $t_{1L},t_{1M},t_{1R}$. Because the error of timing measurement is different and nearly independent in every relative distance measurement, the spectrum of the position error caused by this factor will spread to higher frequency range.
The high frequency component of error plays a huge role in whether the drone can hover stably. When it is much higher, it will introduce a large error in velocity and acceleration estimates and saturates the PID controller, causing the drone to lose control.
Figure 13: Setup for horizontal and vertical movement experiments. (a) Shows the setup for the horizontal movement experiment, and (b) shows the setup for the vertical movement experiment. Blue arrows mean the beacons can communicate with the MP3, and green arrows shows the target movement trajectory. Figure is not drawn to scale.

### v2

@@ Abstract
Communication and position sensing are among the most important capabilities for swarm robots to interact with their peers and perform tasks collaboratively. However, the hardware required to facilitate communication and position sensing is often too complicated, expensive, and bulky to be carried on swarm robots. Here we present Maneuverable Piccolissimo 3 (MP3), a minimalist, single motor drone capable of executing inter-robot communication via infrared light and triangulation-based sensing of relative bearing, distance, and elevation using message arrival time. Thanks to its novel design, MP3 can communicate with peers and localize itself using simple components, keeping its size and mass small and making it inherently safe for human interaction. We present the hardware and software design of MP3 and demonstrate its capability to localize itself, fly stably, and maneuver in the environment using peer-to-peer communication and sensing.
@@ I Introduction
Additionally, a wide variety of tasks, including search and rescue, transportation, or drone shows, can involve direct or indirect interaction with humans. But highly capable multi-motor drones tend to be heavy and unsafe for human interaction without safety equipment [21], a risk that is amplified when interacting with a large number of UAVs. While there are impressive works on active avoidance of obstacles which partially addressed this issue [22, 17], it would be preferred if the drone could be light enough to be inherently safe [23] to interact with.
@@
In this paper, we present a single motor nano [30, 14] aerial vehicle, MP3, with a novel, minimalist peer-to-peer communication and sensing system. MP3 is similar to its predecessors: its only motor is offset from its center of mass to provide maneuvering ability similar to the original Maneuverable Piccolissimo (MP) [27] and its controller is similar to MP2 [15]. However, MP3 is equipped with a novel communication and sensing system as well as a completely new chassis, electrical, and software design. MP3 is capable of communicating with neighboring beacons or other MP3s using infrared light (IR), and sensing its relative bearing, distance, and elevation to neighbors using a single-drone triangulation method. Using this information, MP3 can subsequently determine its global Cartesian position with millimeter accuracy. Using the global position information, MP3 can control its position in 3D space with a single actuator. The simplicity of MP3 allows it to be 20.0g in weight and 112mm in diameter. The small mass and the partially enclosed propeller make MP3 inherently safe for human interaction even with collisions [23]. We present the hardware and software design of MP3, with emphasis on the novel communication and sensing system, and validate its capability with experiments.
Figure 2: An exploded view of MP3. MP3 has 7 major parts: 3D printed main and top chassis, a main PCB, a top transmitter, a motor-propeller assembly, and two batteries. The bottom transmitter is placed at the center of the bottom side of the main PCB.
@@ II-A Overall Concept
MP3, as shown in Fig. 1 and 2, is a minimalist single-rotor drone with on-board communication and position sensing capabilities using a novel optical communication and sensing system. Unlike ordinary drones, MP3 has only one motor offset from its center of mass, two rigid bodies (the rotor fixed to a propeller and the stator fixed to the main chassis), and no inertial measurement unit (IMU). As the motor spins, the body of MP3 spins like a Frisbee in the opposite direction as shown in Fig. 1, and the gyroscopic effect of the spinning body keeps the MP3 stable. By pulsing the thrust at different times, we can tilt the MP3 and control its 3D position in flight [15].
@@
Figure 4: MP3’s communication and sensing system. (a) shows the receiver system (RX) on MP3, (b) shows the top and bottom transmitter system (TX) on MP3, and (c) shows the transmitter system on the Beacon. Note that the transmitter system on the MP3 and on the Beacon are the same.
@@ II-B The Communication and Sensing System
Based on this principle, we designed the communication and sensing system hardware as shown in Fig. 4. Omindirectional transmitters on MP3s blink their infrared (IR) LED to broadcast messages in the form of encoded pulses, and MP3s can receive messages and localize themselves using the three photodiodes on them.
@@ II-B1 Theory of Operation
To illustrate how MP3s achieve relative position sensing and absolute localization, we consider a simplified theoretical model as shown in Fig. 5, where two point-like transmitters are constantly emitting light omnidirectionally and broadcasting their global position $\vec{p}_{1},\vec{p}_{2}$ via IR communication. A simplified MP3 is trying to determine its relative bearing, distance, and elevation to the transmitters as well as its global position. The MP3 is modeled as a solid body rotating at a constant angular velocity $\omega\hat{z}$, and is equipped with three light receivers with planer fields of view (FOV), i.e. a receiver will only pickup light signal when the transmitter is in its FOV. Two light receivers (right and left) are oriented orthogonal to the MP3 horizontal plane, and one light receiver (middle) is oriented at an angle ($\phi$) from orthogonal. As the MP3 rotates, its left, middle, and right receiver will pickup signals from transmitter 1 at time $t_{\text{1L}},t_{\text{1M}},t_{\text{1R}}$ respectively, so we can determine the distance and elevation angle between the MP3 and transmitter 1 as:
@@
We can obtain $r_{2},\alpha_{2}$ from transmitter 2 similarly. Furthermore, at time $t_{1}=(t_{\text{1L}}+t_{\text{1R}})/2$ and $t_{2}=(t_{2L}+t_{2R})/2$, the MP3 is facing transmitter 1 and 2 respectively, so we can express the angle between transmitter 1 and transmitter 2 as
@@
While $\sigma_{\alpha_{1}}$ and $\sigma_{\theta_{12}}$ are on the same order of magnitude as $\omega\sigma_{t}$ and are not very sensitive to relative position between robots, $\sigma_{r_{1}}$ is proportional to $r_{1}^{2}$, meaning that the localization is more accurate when drones are closer to each other. This feature is helpful for collision avoidance in dense swarms.
Unfortunately, non-ideal factors make the implementation of this theory more complicated. These factors include non-ideal sensor optics, limited time resolution constraints by the communications protocols, etc.
@@ II-B3 Electrical Design
MP3 can only receive messages from others when its receivers are facing a transmitting MP3. MP3s typically spin at 25Hz and each of their receivers have a ±$\pm 10^{\circ}$ horizontal FOV, meaning that, in each revolution, a MP3 only has approximately 2-5ms to communicate with another MP3, and in the worst case scenario this gives a communication latency of 40ms. To facilitate communication with multiple MP3s using time division multiple access (TDMA), each MP3 could take no more than 500~1000us to transmit all necessary information to a receiver. If we want to communicate tens of bytes over this period, we need a system capable of communicating at around 1Mbps. This means that we could not use commercially available IR transceivers but had to implement our own.
@@
The receiver circuit is shown in Fig. 8. PCB layout and filter are designed to reduce electromagnetic interference of the motor and filter out almost all environmental light. The analog part of the circuit has a gain of 67mV/uA, an input referred noise as low as 36nA RMS, and a -3dB passband of [0.42MHz,10.5MHz]. The overall circuit can reliably pickup photocurrent as low as 140nA, enabling all manufactured MP3s to communicate across at least 50cm free space with more than 95% success rate using non-directional transmission and reception.
@@ II-B4 Communication Protocol Design
When designing the communication protocol, we want it to be general-purpose, to be simple to implement, to have multiplexing ability, and to provide accurate enough timing information $t_{\text{1L}},t_{\text{1M}},t_{\text{1R}}$ to facilitate position sensing.
To make our protocol general-purpose, we want it to handle messages of different length and content in the same way. This would allow a MP3 to receive communicated positions from other MP3s, as well as other task-dependent messages in future work. A natural approach would be to use packets: a complete message is transmitted in one or multiple 30-bit packets, each containing 2 bytes of data and additional metadata like the ID of the MP3 transmitting the message, the ID of the packet, the cyclic redundancy check (CRC) value of the message, etc. A MP3 can obtain timing information by looking at when it starts to receive packets from another MP3.
@@ II-B5 Global Localization Algorithm
We can now bring everything together. In section II-B1, we have already presented the basic theory to use the time when we received messages from other transmitters to determine relative position. The optical and electrical designs (section II-B2 and section II-B3) make MP3s as close to the theoretical model as possible. Finally, we can use the proposed communication system (section II-B4) to obtain the global position of neighboring MP3s. In this section, we present a simple way of using these relative distance and angle measurements and neighboring MP3s’ global position information to determine MP3’s global position. We execute localization in three steps to simplify the computation without sacrificing much accuracy:
@@
    Determine the global horizontal position $s_{X},s_{Y}$ and the time $t_{x}$ when MP3 is facing $\hat{x}^{+}$ direction using angular velocity of the drone $\omega$, the global horizontal position of the transmitters $(\vec{p}_{i})_{XY}$, the time when MP3 is facing these transmitters $t_{i}$, and the measured relative horizontal distance $r_{i}$.
@@
In the subsequent discussion, we make the following assumptions:
@@
    MP3’s geometric axis is always vertical, and it spins along its geometric axis.
@@
The first assumption will introduce error into vertical position estimation as MP3s’ geometric axis and angular velocity axis will deviate slightly, but usually they are only off by a few degrees and we are willing to sacrifice the precision in vertical position for faster computation. The second and third assumptions are quite accurate as in experiments the angular velocity changes by <0.5% and the position usually changes by <5mm in one revolution. While the measured variables are not strictly Gaussian distributed, the fourth assumption is added for ease of explanation without compromising the core idea.
@@
For the second step, we will determine the most likely $t_{x},s_{X},s_{Y}$ given $\vec{p}_{i},r_{i},t_{i}$ and their uncertainty estimates $\sigma_{r_{i}},\sigma_{t}$.
$\mathop{\mathrm{argmin}}\limits_{t_{x},s_{X},s_{Y}}\sum\limits_{i=1}^{N}\left(\frac{\Delta_{X}^{2}+\Delta_{Y}^{2}}{\sigma_{r_{i}}^{2}}\right.\left.+\frac{\mathop{\mathrm{atan}}(\Delta_{X},\Delta_{Y})-\omega(t_{i}-t_{x})}{\omega^{2}\sigma_{t}^{2}}\right)$ (8)
@@
which is the root mean square (RMS) of uncertainty in X and Y direction. For simplicity of computation, in the third step we will consider $s_{X},s_{Y}$ as independent random variables that both have uncertainty of $\sigma_{XY}$.
@@
where $\delta_{i}$ is the uncertainty of $(\vec{p}_{i})_{Z}-\sqrt{\Delta_{X}^{2}+\Delta_{Y}^{2}}\tan\alpha_{i}$
@@
In experiments, we additionally applied an exponential filter with time constant of 0.06s to the obtained position ${s_{X},s_{Y},s_{Z}}$ to reduce the noise of position estimates.
@@ II-C The Mechanical System
Several other changes include: a brushless motor instead of a brushed motor to improve longevity and battery life, carbon fiber-infused nylon for chassis to reduce weight and increase strength, and wings added to the side of MP3 to improve passive stability and reduce rotation rates.
@@ III Drone Flight Experiments
For each experiment, we collected the reference trajectory of the MP3s using two cameras, one over the apparatus and one on the side, and we also recorded the estimated trajectory of the MP3 which is computed on-board using only relative measurements and data communicated through our IR communication system. Data are recorded after MP3 is flying steadily, about 10~20s after launch.
@@ III-A Position Holding Experiment
In the position holding experiments, MP3 localizes itself by sensing three neighboring beacons. MP3 then tries to remain at a predefined set point at the center of the three beacons. The setup is shown in Fig. 10. We plot the reference trajectories of MP3 in all flights in Fig. 11. The trajectories indicate successful and stable position holding, since MP3 is staying within 30mm around the origin most of the time (in contrast, the radius of MP3 is 56mm). The root mean square error (RMSE) of the drone’s reference position is {17.6, 22.5, 12.7}mm in X, Y and Z direction respectively.
@@
Figure 12: Distribution and spectrum of localization error of MP3 in position holding experiments. (a) shows the localization error’s distribution, and (b) shows the localization error’s spectrum.
A way to evaluate the precision of localization is to calculate the error between the robot-sensed trajectory and reference trajectory. Fig. 12(a) shows the distribution of the localization error in X, Y and Z direction, which resemble normal distribution with $\sigma=\{2.4,2.0,3.6\}$mm respectively. This suggests that our sensing system and localization algorithm can achieve millimeter level accuracy. The distribution’s offset in X and Y direction may be due to inaccurate calibration of the camera over the entire flight arena or slight displacement of the beacons, and the offset in Z direction may also be caused by the MP3’s tilt.
Besides the localization error’s distribution, we are also interested in the spectrum of error, as the higher frequency components in error will be amplified more when we take derivatives to compute the velocity and acceleration. We split the obtained trajectory data into chunks of 5 seconds and resample them at 100Hz. We then compute the discrete Fourier transform (DFT) of the error samples and plot the RMS density of error signal at different frequencies in Fig. 12(b). The lower frequency (<2Hz) components are dominated by systematic error like the inaccurate placement of beacons or inaccuracy in the functions that converts time $t_{1L},t_{1M},t_{1R}$ to distance and elevation angle. These sources are dependent on the position of the drone which is slowly changing. The higher frequency (>2Hz) components, however, mainly come from the discrete and random nature of timing measurements $t_{1L},t_{1M},t_{1R}$. Because the error of timing measurement is different and nearly independent in every relative distance measurement, the spectrum of the position error caused by this factor will spread to higher frequency range.
The high-frequency error component significantly effects the drone’s stability during hovering. When it is much higher, it will introduce a large error in velocity and acceleration estimates and saturate the PID controller, causing the drone to lose control.
Figure 13: Setup for horizontal and vertical movement experiments. (a) shows the setup for the horizontal movement experiment, and (b) shows the setup for the vertical movement experiment. Blue arrows mean the beacons can communicate with the MP3, and green arrows show the target movement trajectory. Figure is not drawn to scale.